OpenStudy (anonymous):

What are the explicit equation and domain for an arithmetic sequence with a first term of 5 and a second term of 2? @jim_thompson5910 @myininaya @mathstudent55 @SolomonZelman @zepdrix @dan815 @wio @abb0t @iambatman

jimthompson5910 (jim_thompson5910):

hint: an = a1 + d(n-1) an is the nth term a1 is the first term d is the common difference n is the term number

OpenStudy (anonymous):

so, an=5+3(n-1)

jimthompson5910 (jim_thompson5910):

first term: a1 = 5 second term: a2 = 2

jimthompson5910 (jim_thompson5910):

common difference: (second term) - (first term) = 2 - 5 = -3 so d = -3 and NOT d = 3

OpenStudy (anonymous):

ohhh! So an=5-3(n-1)

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

But is it integers when n>=1, or n>=0?

jimthompson5910 (jim_thompson5910):

n starts at n = 1, so n >= 1

jimthompson5910 (jim_thompson5910):

it's easiest to start at n = 1 to refer to the first term n = 0 referring to the first term confuses things a bit

OpenStudy (anonymous):

Thank you!!!!

jimthompson5910 (jim_thompson5910):

np